OpenStudy (anonymous):

Help! I will medal!! Given the function f(x) = 2x, find the value of f−1(32). f−1(32) = 0 f−1(32) = 1 f−1(32) = 5 f−1(32) = 16 Can someone go through the steps for me!!

OpenStudy (anonymous):

f−1(32) = 16. Because, 2^x=32 Then you make them both have the same base, so we then have 2^x=2^16. We can then eliminate the bases so, x=16. f^-1(x)=16 f^-1(32)=16